. A diesel engine uses 54.5 kg of fuel oil per hour of calorific value 45 MJ/kg. The thermal efficiency of the engine is 25 per cent. Determine the power output of the engine in kilowatts
1
Expert's answer
2020-12-14T12:14:36-0500

Input fuel to engine is 54.5kg/hour

and input power to engine is "54.5\times 45 \times 10^6" joules/hour

efficiency of engine is 25%

Hence output power of engine is "\cfrac{2452.5}{4}\times 10^6" Joules/hour

In kilowatts, output power is "\cfrac{613.1}{3600}\times 10^3" Kilo-joules/sec

or 170.3 kilowatts.